OpenStudy (anonymous):

A truck has a full 50-gallon gas tank. It uses 7 gallons on the first part of its journey, 13 gallons on the second part of its journey, and 15 gallons on the third part of its journey. How many gallons of gas remain in the gas tank?

OpenStudy (anonymous):

7 + 13 + 15 = x 50 - x = answer

OpenStudy (anonymous):

anotherwords, 15 gallons

OpenStudy (anonymous):

subtract the amount of gas used from the original 50 gallons 50 - 7 -13 - 15 = 15 gallons remaining